Write an equation of the line in standard form with the given slope through the given point.
Slope = 4, (-2, -5)
I know standard form is ax+by=c but I am not sure where all the numbers go. This is what I have:
4(-2) + b(-5) = c
I feel like it's not right. Help please?

Answer by Edwin McCravy(20086)

slope 4, through the point (-2, -5)

First we use the point-slope form:

Point-slope formula:

y - y1 = m(x - x1)

where m = 4 and (x1,y1) = (-2,-5)

y - (-5) = 4(x - (-2))

   y + 5 = 4(x + 2)
   y + 5 = 4x + 8
       y = 4x + 3

Now we must get it in the form ax + by = c

       y = 4x + 3

Add -4x to both sides

 -4x + y = 3

But the coefficient of x must be positive to have the standard form,
so we multiply through each term by -1, which changes all the signs:

 4x - y = -3

Edwin